OpenStudy (anonymous):

x-4/x+2<1 because the x-4 is divided by x+2 well times both sides by x+2 so we now have the following (x+2) (x-4)<1(x+2) than we distr (x+2) (x-4) which we get x^2-3<1 did i miss something here plz help

OpenStudy (anonymous):

i am here

OpenStudy (anonymous):

am I on the right track?

OpenStudy (campbell_st):

you would normally multiply by the square of the denominator \[(x + 2)(x -4) < (x + 2)^2\] because if you multiply by (x + 2) you have \[\frac{(x - 4)}{(x + 2)} \times (x + 2) < 1 \times (x + 2)\] which becomes \[\frac{(x -4)(x+2)}{(x +2)} < (x +2) \] cancel the common factor in the fraction leaves \[(x -4) < (x + 2)\] hope this helps

OpenStudy (anonymous):

where is the 0ne?

OpenStudy (campbell_st):

well 1 times anything is 1 so 1 times (x + 2) is x + 2